A study utilizes ANOVA to examine the average distribution rates of an antibiotic in infants, toddlers and school age children. The results include an F statistic that is 1.04. You know this means:
Answer #1

This means that the result is not significant at 0.01, 0.05 and 0.10 significance levels.

p-value for F =1.04 and at df(numerator) =number of factors - 1 =3 - 1 =2 and df(denominator) =overall sample size - number of factors =N - 3 (N is not given) is around 0.3

So, p-value of about 0.3 >0.01; >0.05 and >0.10 significance levels. Hence, we failed to reject the null hypothesis(H0) that says "no two population means differ significantly".

Thus, there is no sufficient evidence to claim that at least two population means differ significantly(H1: alternative hypothesis).